OpenStudy (anonymous):

Y=-X^2+14X-49 Find the -X INTERCEPT and the coordinates of the vertex of the parabola

OpenStudy (lgbasallote):

x intercept is the value of x when y = 0..so set y to zero then solve for x...the vertices i forgot @dpaInc may know though

OpenStudy (anonymous):

for the vertice, get it in this form by completing the square: 4a(y-k) = (x-h)^2 where (h, k) is your vertex.

OpenStudy (anonymous):

there's actually another way of getting the vertex but I can't recall it. it went something like b/2a ?... sound familiar?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

??

OpenStudy (lgbasallote):

sounds like quadratic formula @dpaInc =))

OpenStudy (anonymous):

The x-coordinate for the vertex of the parabola is x=-b/2a

OpenStudy (anonymous):

that's it....

OpenStudy (anonymous):

And then plug in to solve for the y-coordinate.

OpenStudy (anonymous):

where dhe go?

OpenStudy (anonymous):

I think he got it. :D

OpenStudy (dumbcow):

factor and set equal to 0 --> -(x-7)^2 = 0 x = 7

OpenStudy (anonymous):

also i think i have a formula for y coord of vertex from a, b, c y = c - (b^2)/4a

OpenStudy (anonymous):

it sounds like it's gonna be way easier if you just chug the value of the x-coordinate into the equation. :)

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

just threw that in there for the sake of it

OpenStudy (anonymous):

but that's a good answer still..

OpenStudy (anonymous):

Notice that \[y= -x^2+14x-49 =- (x^2 -14 x +49)=-(x-7)^2 \] So the x-intercept is x=7 and the vertex is (-7,0)